# SDFs from Unoriented Point Clouds using Neural Variational Heat Distances

### v2

@@ 4.4 Application: Geometric Queries
Figure˜15 shows an example result. Note that in general, neither $\min\{\phi_{1},\phi_{2}\}$ nor $\max\{\phi_{1},\phi_{2}\}$ is an SDF [22].
@@ 5 Limitations
Our method has a few limitations. To ensure that we solve only first-order variational problems that are convex with respect to the function gradients, we train two separate networks, one for the heat diffusion time step $u^{\tau}$, and one for the SDF $\phi$. We also observe a moderate loss of detail, which is a common drawback of a neural representations. Additionally, singularities such as crease lines lead to locally larger errors in the SDF. The proposed method for computing the sets $\mathcal{B}^{\pm}$ does not generalize to shapes with interior voids or disconnected internal structures, albeit such cases are relatively uncommon. For these challenging cases, our method can be extended by using Dipole Normal Propagation (DNP) [24] for point cloud orientation and then generalized winding numbers [18] to determine $\mathcal{B}^{\pm}$. In Figure˜16, we show a failure case of our box algorithm, while the extended pipeline succeeds—at the cost of increased computation time.
@@ Acknowledgments
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 211504053 – Collaborative Research Center 1060 and via Germany’s Excellence Strategy project 390685813 – Hausdorff Center for Mathematics. Furthermore, this project has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 101034255, as well as the Israel Science Foundation (grant No. 1073/21).
@@ Appendix A Proofs
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|+|A(\phi_{j},n_{k}-n)|+|A(\phi,n_{k}-n)|$
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|+(\|\nabla\phi_{j}\|_{L^{2}}+\|\nabla\phi\|_{L^{2}})\|n_{k}-n\|_{L^{2}}$
@@ Appendix B Additional Experimental Details
Figure 17: Marching Cubes result for zero-level sets of our neural SDF. Models used in the quantitative evaluation in Figure˜11.
@@
Here, we provide some additional qualitative and quantitative evaluation and details on the used data. In Table˜3 we provide quantitative evaluation for the noisy and sparse input point cloud of the capped torus shape from Figure˜13.
In Figure˜17, we show Marching Cubes results for the zero-level sets of our neural SDF on the set of shapes used in the quantitative evaluation in Figure˜11.
In Figure˜18, we show the zero-level set results for some additional shapes. They were all computed with the same parameter set that is mentioned in the main text. Notably, no parameter tuning is required to reconstruct thin structures like the birds claws.

### v3

@@ 4.4 Application: Geometric Queries
Figure˜16 shows an example result. Note that in general, neither $\min\{\phi_{1},\phi_{2}\}$ nor $\max\{\phi_{1},\phi_{2}\}$ is an SDF [22].
@@ 5 Limitations
Our method has a few limitations. To ensure that we solve only first-order variational problems that are convex with respect to the function gradients, we train two separate networks, one for the heat diffusion time step $u^{\tau}$, and one for the SDF $\phi$. We also observe a moderate loss of detail, which is a common drawback of a neural representations. Additionally, singularities such as crease lines lead to locally larger errors in the SDF. The proposed method for computing the sets $\mathcal{B}^{\pm}$ does not generalize to shapes with interior voids or disconnected internal structures, albeit such cases are relatively uncommon. For these challenging cases, our method can be extended by using Dipole Normal Propagation (DNP) [24] for point cloud orientation and then generalized winding numbers [18] to determine $\mathcal{B}^{\pm}$. In Figure˜17, we show a failure case of our box algorithm, while the extended pipeline succeeds—at the cost of increased computation time.
@@ Acknowledgments
This work was supported by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) via project 211504053 – Collaborative Research Center 1060, project 539309657 – Collaborative Research Center 1720, and via Germany’s Excellence Strategy project 390685813 – Hausdorff Center for Mathematics. Furthermore, this project has received funding from the European Union’s Horizon 2020 research and innovation program under the Marie Skłodowska-Curie grant agreement No 101034255, as well as the Israel Science Foundation (grant No. 1073/21).
@@ Appendix A Proofs
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|$
$\qquad+|A(\phi_{j},n_{k}-n)|+|A(\phi,n_{k}-n)|$
$\leq|A(\phi_{j},n_{k})-A(\phi,n_{k})|$
$\qquad+(\|\nabla\phi_{j}\|_{L^{2}}+\|\nabla\phi\|_{L^{2}})\|n_{k}-n\|_{L^{2}}$
@@ Appendix B Additional Experimental Details
Here, we provide some additional qualitative and quantitative evaluation and details on the used data. In Table˜3 we provide quantitative evaluation for the noisy and sparse input point cloud of the capped torus shape from Figure˜12.
In Figure˜18, we show Marching Cubes results for the zero-level sets of our neural SDF on the set of shapes used in the quantitative evaluation in Figure˜10.
In Figure˜19, we show the zero-level set results for some additional shapes. They were all computed with the same parameter set that is mentioned in the main text. Notably, no parameter tuning is required to reconstruct thin structures like the birds claws.
Figure 18: Marching Cubes result for zero-level sets of our neural SDF. Models used in the quantitative evaluation in Figure˜10.